Direct Materials Variances

I-Time, Inc., produces electronic timepieces. The company uses mini-LCD displays for its products. Each timepiece uses one display. The company produced 480 timepieces during October. However, due to LCD defects, the company actually used 500 LCD displays during October. Each display has a standard cost of $6.8. The company purchased 500 LCD displays for October production at a cost of $2,805.

Determine the price variance, quantity variance, and total direct materials cost variance for October.

Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Price variance $ Favorable
Quantity variance $ Unfavorable
Total direct materials cost variance $ Favorable
Answer #1

Price Variance = (SP-AP) * AQ

= (6.8 - 2805/500) * 500

= - 595 Favourable (minus sign here)

Quantity Variance = (SQ-AQ) * SP

= (480 - 500) * 6.8

= 136 Unfavorable

Cost variance = 595F + 136U

= - 459 Favourable (minus sign here)